OpenStudy (music0203):

solve the equation using the Zero-Product Property. -2x(5x-2)=0

OpenStudy (kidrah69):

5x-2 = 0 -2x=0

OpenStudy (kidrah69):

Now you just solve it :)

OpenStudy (jjuden):

5x-2 = 0 -2x=0

OpenStudy (kidrah69):

@jjuden thats what i wrote :)

OpenStudy (music0203):

wait, but how do you solve it from there?

OpenStudy (jjuden):

@kidrah69 i know some people need 2 answers to make sure though :)

OpenStudy (kidrah69):

Just like how u would solve any equation

OpenStudy (kidrah69):

isolate X

OpenStudy (kidrah69):

@jjuden oh ok :)

OpenStudy (kidrah69):

5x-2 = 0 add 2 to both sides 5x=2 then divide x=5/2

OpenStudy (kidrah69):

Try the 2nd one out -2x=0

OpenStudy (music0203):

divide -2 by both sides?

OpenStudy (kidrah69):

yep! :)

OpenStudy (music0203):

so thats just 0. now what?

OpenStudy (kidrah69):

thats the answer silly!

OpenStudy (kidrah69):

You will have 2 answers x=5/2 x=0

OpenStudy (music0203):

the answer choices are A. 0,2/5 B. 0,2 C. 0,-2/5 D.0,-2

OpenStudy (music0203):

so its 2/5 not 5/2....

OpenStudy (kidrah69):

right sorry >_< made a typo :( yeah 2/5

OpenStudy (music0203):

so its A?

OpenStudy (kidrah69):

yep =)

OpenStudy (music0203):

thank you! do you want a metal?

OpenStudy (kidrah69):

it does not matter to me :) you can give one at your own will :) im glad i can help =) you are welcome!